A particle of mass 7.5 kg is moving at a steady speed of 12ms-1 when it meets with a constant resistance of 10N. How long does it take to come to rest?
1
Expert's answer
2020-03-31T09:26:50-0400

The acceleration of the particle

"a=\frac{F}{m}=\frac{10\:\rm N}{7.5\:\rm kg}=1.3\:\rm m/s^2."

The time of motion to the rest

"t=\frac{v}{a}=\frac{12\:\rm m/s}{1/3\:\rm m/s^2}=9\:\rm s."